OpenStudy (anonymous):

how do you solve this? Do you think the equation x+2=x-3 has more than one real number solution? If so, how many solutions does it have? explain you reasoning. ~this is a study question on my midterm coming up.... the answer is no, but how do you work the problem?

OpenStudy (anonymous):

the answer is no because whatever \(x\) is on the left, it is the same number on the right. you cannot add 2 to a number, subtract 3 from that same number, and end up equal

OpenStudy (anonymous):

in other words, this equation cannot have any solution at all

OpenStudy (anonymous):

You can verify by solving using standard procedures \( x+2=x-3\) subtract 2 from both sides \( x=x-5\) subtract x from both sides \( 0=-5\) That statement is never true, so no value for x can solve it.